𝛼 = 0.01; two-tailed test; 𝑛 = 15; 𝜎 = 5 .


1
Expert's answer
2021-07-07T10:32:18-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=\mu_0"

"H_0:\mu\not=\mu_0"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.01," and the critical value for a two-tailed test is "z_c=2.5758."

The rejection region for this two-tailed test is "R=\{z:|z|>2.5758\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu_0}{\sigma/\sqrt{n}}"

If it is observed that "|z|>z_c," it is then concluded that the null hypothesis is rejected. There is enough evidence to claim that the population "\mu" is different than "\mu_0," at the "\alpha=0.01" significance level.


If it is observed that "|z|\leq z_c," it is then concluded that the null hypothesis is not rejected. There is not enough evidence to claim that the population "\mu"  is different than "\mu_0," at the "\alpha=0.01" significance level.